Please help with creating journal entries for the following

ABC Corp.

Balance Sheet

As of December 31, 2006

Assets

Liability

Cash

$6,000

Accounts Payable

$3,000

Accounts Receivable

$4,000

Stockholders' Equity

Common Stock

$3,000

Additional Paid-in Capital

$2,000

Retained Earnings

$2,000

Total Liability and Stockholders' Equity

Total Assets

$10,000

$10,000

Below is a summary of events experienced by ABC Corp. during 2007

  1. On Jan 1st, 2007, ABC Corp. purchased equipments for $800. The equipment has a useful life of four years and ABC Corp. use straight-line depreciation.
  2. On April 1st, 2007, ABC purchased a one-year insurance policy for $1200. The policy is effective immediately.
  3. ABC Corp. declared and paid cash dividends of $1,000.
Answer #1
Date Particulars Debit Credit
A 01 January 2007 Equipments                          800.00
Cash                            800.00
B 01 April 2007 Insurance expense                      1,200.00
Cash                        1,200.00
C 31 December 2007 Retained earnings                      1,000.00
Dividends payable                        1,000.00
(journal entry for dividends declared)
D 31 December 2007 Dividends payable                      1,000.00
Cash                        1,000.00
(journal entry for dividends paid)
E 31 December 2007 Depreciation                          200.00
Accumulated depreciation                            200.00
($800/4 years = $200 per year)
F 31 December 2007 Prepaid insurance                          300.00
Insurance expense                            300.00
(adjusting entry)

Explanations: The insurance amount of $1200 is for 12 months and so monthly expenses = 1200/12 = $100 per month.

Expenses that should be recorded in books, as per accrual accounting system, should be =  expenses for the period 1st April to 31st December i.e. a period of nine months. The amount will be = 9 months*$100 per month = $900. The remaining amount of 1200-900 = 300 will be recognized as prepaid insurance expense and will stand as a current asset in balance sheet.